OpenStudy (anonymous):

write an equation in the form y =mx+b for the line that passes through the point (-5,-1) and has a slope of -2

OpenStudy (tkhunny):

Well, do it. You have "y = mx + b" and you are give the slope. Simply substitute that. y = -2x + b Now what? That point was given for a reason.

OpenStudy (anonymous):

do i subtract -2 from -1

OpenStudy (tkhunny):

Why would you do that? Use the given point. (x,y) = (-5,-1)

OpenStudy (anonymous):

i dont understand

OpenStudy (tkhunny):

What does (-5,-1) mean when it indicates a point on a line?

pooja195 (pooja195):

Before any of the we need to do the point slope formula \[\huge~y-y_1=m(x-x_1)\] m=slope (x1,y1)

OpenStudy (anonymous):

y--2=m(x-x1)

pooja195 (pooja195):

We also have the slope :) m=-2 try again

OpenStudy (tkhunny):

Yes, we can use the point-slope formula, but then you have to remember that. I prefer simply thinking. Please answer my question. What does (-5,-1) mean when it indicates a point on a line?

OpenStudy (anonymous):

y--1=m(x-x1

pooja195 (pooja195):

Still wrong. Let me help you out :) -5 = x1 -1 = y1 -2=m

pooja195 (pooja195):

Knowing that can you plug in the numbers?

pooja195 (pooja195):

Stuck?

OpenStudy (tkhunny):

 "plug in numbers" doesn't mean anything. The idea of "substitution" is universal and should be burned into your brain.